OpenStudy (anonymous):

What is the equation in standard form of the line through (4,1) and (-2,3) help!!!

OpenStudy (anonymous):

x - 3y = 1

OpenStudy (anonymous):

thnx but can u plz tell me how u got that?

OpenStudy (anonymous):

First, find the gradient using: \[m = (y_{2} - y_{1})/(x_{2} - x_{1})\]\[m = (3 - 1) / (-2 -4)\]\[m = 1/3\] Using the formula for finding eqn of a line: \[y - y_{1} = m(x - x_{1})\]\[y - 1 = 1(x - 4)/3\]\[y = 1(x)/3 - 1/3\] Expressing in standard form: \[x - 3y = 1\]

OpenStudy (anonymous):

I am new to this site so pardon my equations. Learning how to use the equation editor

OpenStudy (anonymous):

ooo thank you!!!!!!!

OpenStudy (anonymous):

You welcome :)